NAOJ GW Elog Logbook 3.2

To access to the MC finesse, we tried to use the following fit : T0/ ( 1 + 4R / (1-R)² * sin² ( pi (x-x0)/FSR ) ) + bg
Attached to this report are two results we could get : one without normalization, one with normalization.
By using Finesse = 4R / (1-R)², we could get the following results :
First case : 209
Second case (with normalization) : 360
By checking the ratio FSR over peak width : Finesse = 150.
We will try to undestand better how to fit this function

Today, it was not possible to lock the cavity because of a drift of the end mirror.
When we went to the end room to reset the local control, the temperature and humidity of the end room was quite high.
It appears that a air ventilation system was off. As soon as we turned it on, it started to feel more confortable.
After resetting the end mirror local control, it was possible to lock the cavity again.

I installed the 1310nm laser and the relative Imaging Unit
items:
- Laser controller
- fiber output
- golden half-inch mirror
- golden small prism mirror (before the cross point)
On the Imaging unit translation stage:
- golden large prism mirror (after the cross point)
- XY lens mount
- coated half ball
- Photo Detector

The Spectrum Analyzer provided data in ".DAT" file.
By using a program provided by Tatsumi-san, we were able to convert these data in ".DOT" file and then use Matlab to plot them.
This ".DOT" file is divided in 3 columns : frequency, magnitude and phase of the transfer function.
Attached to this entry are the open-loop, the electronic-loop and the optical loop.
They seem to be coherent with what the spectrum analyzer displayed during the measurement.

Before install everything, we checked the beam size after the transmission of the beam splitter, which is the same at the reflection. Take the origin at the front surface of the beam splitter, the beam waist is at -9.72cm, size is 53.36um. Then the first lens we use is 100mm, 5cm from the origin, but at the focal plane the beam diameter is larger than 1.3mm, which is the maximum the AOM can work according to the data sheet. So the AOM was put around 11cm, connected with the power, the AOM shows very clear diffraction orders.
Then another 100mm lens was put at 25cm. After putting this lens, we did some measurement of the beam size, and found the position to put the third lens in order to have a good size at the 2inch mirror of the telescope(less than 1.3mm). The third lens is 175mm, was put at 48.5cm. Then we checked the beam shape far, there is no obvious astigmatism as far as we can see. Measuring the beam again and we got the result that the beam at 2 inch mirror should be less than 1.2mm.
But then we found out if we want to change the MZ design and put another beam splitter after the one we have now, the first 100mm lens should be put further to give enough space to the BS. With Jammt, I did the simulation and found out if we move everything together 2.5cm further from the position mentioned before, the 175mm lens was not capable to focus the beam enough on the 2inch mirror. I tried other focal length, the 200mm should work.
The other problem is after moving(pic 1), the beam after the first lens will diverge more compared to the previous design(pic 2),sSince the two green dash lines in two pictures have the same distance. Divergence of the beam is one of the question we concern most, so I checked on the bench, after changing the AOM still works well.
It seems this new configuration is acceptable, we are going to install another beam splitter first and start to change the lenses.

The amplitude of the loop transfer functions plotted so far are actually the square of the real amplitude. The problem comes from the way I treated data saved by the spectrum analyzer. Each file is composed of 3 columns: frequency, real part (a) and imaginary part (b) of the TF. Of course amplitude and phase are recovered by doing:
Amplitude = sqrt (a^2 +b^2)
Phase = angle (a+i*b)
Due to an oversight, I had replaced the sqare root with the absolute value in the amplitude computation. This explain the unexpected behaviour (1/f^2 instead of 1/f) of the openloop TF around the UGF.
We will upload soon new TFs measurements (taken by Yuefan and Marc on monday night) properly plotted.

Mode cleaner cavity consists of three mirrors, two of them are flat and one is curve. After the curve mirror there is the PZT to change the cavity length for finding a good mode matching. This piezo is connected to the output of PZT driver whose input is connected to a function generator to provide the scan signal.
Since we only want to do a simple test, so we did not use the telescope design but only one 200mm lens after the beam splitter, then two steering mirrors used to align the cavity, at the output of the MC, a PD with DC output is used to see the modes. The whole configuration shows in pic 1.
We used 25Hz ramp wave with amplitude of 1Vpp to scan the cavity. At the beginning, we only saw some fluctuation but no peaks. When we tried to make the output beam go straight, we were not able to do it.(Always cut by the mirror mount) So we removed the MC and aligned better from the lens, sent the beam after the mirror far enough to make sure it goes along the holes of table. When we put back the MC, we could see some higher modes at the output and also the curve mirror has some transmission beam this time. Put back the PD, we saw pic 2 on the oscilloscope. By checking the beam shape with the curve mirror transmission and the spectrum, we got better mode matching. In pic 3, the highest peak is TEM00, we also checked it by moving the voltage of the PZT driver by hand. I think this means the mechanical part of the MC works well, we are able to align the cavity with this design.

A small problem was encounter : While screwing the cover part of the hole of cavity to the mode cleaner, a screw stayed stuck.
This was due to the fact that too long screw were proposed on the design. Instead of using M4*12, we used M4*8 screw for this part. We also used ethanol to avoid to stuck another screw.
Also, we couldn't find M3*16 screws so we used M3*15 screws.
To protect the wire used for the piezo power supply, we used a small piece of aluminum folded. One part is screwed to the optical bench while the other part hold a adaptor between BNC and the 2 wires for the piezo.

The table is to be updated with the values of yesterday (in blue).
low probe power | high probe power | |||
signal | AC | 23mV | AC | 230mV→295mV |
DC | 440mV | DC | 5400mV→5200mV | |
AC/DC | 0.052 | AC/DC | 0.043→0.057 | |
noise With sample |
AC_rms | 0.1mV | AC_rms | 2.8mV→1mV |
DC | 440mV | DC | 5400mV→5200mV | |
AC_rms/DC | 2.30E-04 | AC_rms/DC | 4.5e-4→1.9e-4 | |
ppm | 884ppm | ppm | 2000ppm→667ppm | |
noise Without sample |
AC_rms | 3μV | AC_rms | 70μV→ 200μV |
DC | 0.65V | DC | 6.5V | |
AC_rms/DC | 4.60E-06 | AC_rms/DC | 1e-5→3e-5 | |
ppm | 18ppm | ppm | 46ppm→108ppm |

This week, in order to check the AOM characteristics, we install the AOM after a beam splitter on the green path. By using a beam splitter before the AOM and 2 powermeters ( 1 one reflexion, the other on the transmission at the output of the AOM ) and checking their ratio, we were able to characterize the AOM despite still having power fluctuations on the green beam. The optical setup used is described in an attached figure.
By changing the RF power send to the AOM, we were able to characterize the AOM 1st order with the use of a gaussian fit ( even if this wasn’t really a gaussian, it helped to locate the maximum) as following :
- Maximum efficiency : 73 % @ RF Power 28.4 dBm ( 692 mW)
The AOM test sheet said that we could expect a 1st order efficiency superior than 85% at 633 nm. In this case, our alignment was approximative as we wanted to check only the response of the AOM to different RF power.
Then we tried to put the AOM on the right position on the optical bench. As the AOM need a small input beam size, we put it in the middle of 2 lenses ( f = 100 mm ) .
At that position, we couldn't see anymore any diffraction order.
First, we checked the green Power Density sent to the AOM. We measure 10W/mm² when the AOM test sheet limit this power density to 2.5 W/mm². Hopefully, we reduced quickly (after few minutes) the laser power down to 2 W/mm². In regard to this, we contact AA Opto-Electronic, manufacturer of this AOM. Following their advice, we check that the crystal was still transparent without any visible damages.
Then, we tried to put the AOM back on the characterization position. We were able to see again diffraction orders. We realize again the characterization of the 1st order efficiency and obtain :
- Maximum efficiency : 69 % @ RF Power 28.3 dBm ( 676 mW) We expect that the difference might be due to misalignment.
After that, we checked the polarization of the green beam using a PBS because this AOM needs a vertical polarization to work. We found that in both positions the green beam has a vertical polarization as we expect.
The last difference is the divergence of the beam. Indeed the beam is really more diverging in the right position (5.6 mrad) than on the characterization position (1.6 mrad) compared to the diffraction angle (16.6 mrad).
To correct this problem we will try to change the lenses configuration in order to obtain a smaller divergence of the beam on the right AOM position.

case | std(AC) | DC | std(AC/DC) | AC | std(AC) | AC/DC | std(AC/DC) | std(ppm) | std(ppm) |
pump | OFF | OFF/ON | OFF | ON | ON | ON | ON | OFF | ON |
without IR filter | 3.4uV | 0.77V | 4.4uV | 0.042V | 128uV | 0.0545 | 1.67E-04 | 16 | 621 |
with IR filter | 2.7uV | 0.45V | 6.1uV | 0.024V | 99.5uV | 0.053 | 2.26E-04 | 23 | 873 |
We found out that the DC changes from 0.46V to 0.44V when switching off the pump. This happens only when there is the sample, this means that some pump is scattered from the sample.
low probe power | high probe power | |||
signal | AC | 23mV | AC | 230mV |
DC | 440mV | DC | 5400mV | |
AC/DC | 0.052 | AC/DC | 0.043 | |
noise With sample |
AC_rms | 0.1mV | AC_rms | 2.8mV |
DC | 440mV | DC | 5400mV | |
AC_rms/DC | 2.30E-04 | AC_rms/DC | 4.50E-04 | |
ppm | 884ppm | ppm | 2000ppm | |
noise Without sample |
AC_rms | 3μV | AC_rms | 70μV |
DC | 0.65V | DC | 6.5V | |
AC_rms/DC | 4.60E-06 | AC_rms/DC | 1.00E-05 | |
ppm | 18ppm | ppm | 46ppm |
Today we checked again the signal level at the above conditions and we found almost the same values of the table above but the AC noise with low probe power and without sample was higher: around 200 μV instead of 70μV.
Then Kuroki suggested to cover the Imaging Unit to protect from wind (as it was in the original setup last year) and the noise became between 50-100μV , then we removed the cover and the noise remained on the same level 50-100μV. We think we should cover better the optical parts, in order to avoid temperature fluctuations which might affect the noise.
The table is to be updated with the values of yesterday (in blue).
low probe power | high probe power | |||
signal | AC | 23mV | AC | 230mV→295mV |
DC | 440mV | DC | 5400mV→5200mV | |
AC/DC | 0.052 | AC/DC | 0.043→0.057 | |
noise With sample |
AC_rms | 0.1mV | AC_rms | 2.8mV→1mV |
DC | 440mV | DC | 5400mV→5200mV | |
AC_rms/DC | 2.30E-04 | AC_rms/DC | 4.5e-4→1.9e-4 | |
ppm | 884ppm | ppm | 2000ppm→667ppm | |
noise Without sample |
AC_rms | 3μV | AC_rms | 70μV→ 200μV |
DC | 0.65V | DC | 6.5V | |
AC_rms/DC | 4.60E-06 | AC_rms/DC | 1e-5→3e-5 | |
ppm | 18ppm | ppm | 46ppm→108ppm |

We tried to calibrate the AOM again. Since the PD has too much effect on the power fluctuation, we decided to put two power meter in 0 and 1st order, so when we change the RF power, we can see the difference between these two orders. But we cannot find a good position to put the power meter that two order is separated enough and also the beam size is smaller than the aperture of the power meter. So we put another beam splitter after the MZ BS, and two power meter on two path of this beam splitter. Then with spectrum analyzer, we did the ratio between these two power in time domain to see the real change effected by the RF power.

In the past days we tried to characterize the locking loop of the filter.
The loop transfer function for the filter cavity (sketched in figure1) is compose by different blocks
- G1 [Hz/V] = piezo actuator
- G2 [Hz/Hz] = SHG
- G3 [Hz/W] = cavity
- G4 [W/V] = photodiode + demodulation
- H [V/V]= servo
- inject noise on perturb
- measure EPS1/EPS2
- inject noise on perturb
- measure piezo mon/EPS2
- inject noise on RAMP
- measure EPS2/piezo mon
The amplitude of the loop transfer functions plotted so far are actually the square of the real amplitude. The problem comes from the way I treated data saved by the spectrum analyzer. Each file is composed of 3 columns: frequency, real part (a) and imaginary part (b) of the TF. Of course amplitude and phase are recovered by doing:
Amplitude = sqrt (a^2 +b^2)
Phase = angle (a+i*b)
Due to an oversight, I had replaced the sqare root with the absolute value in the amplitude computation. This explain the unexpected behaviour (1/f^2 instead of 1/f) of the openloop TF around the UGF.
We will upload soon new TFs measurements (taken by Yuefan and Marc on monday night) properly plotted.

The activity of locking characterization of the past days pointed out some issues that are worth to be reported.
open loop (urad) | closed loop (urad) | |
PR yaw | 1-2 | 0.5-1 |
PR pitch | 4-8 | 2-3 |
BS yaw | 3-4 | 2-3 |
BS pitch | 5-6 | 4-5 |
INPUT yaw | 2-3 | 1-2 |
INPUT pitch | 4-5 | 3-4 |
END yaw | 2-3 | 1-2 |
END pitch | 4-5 | 3-4 |

After a long alignment work we were able to make the IR and the green beam flashing in the cavity at the same time. In this video are shown superposed flashes in trasmission both of the green and the IR (the green beam as been cut at the second 5).
In this configuration we were able to lock the cavity on the green TEM00 but we coudn't check the IR condition since there was too much green light transmitted by the dichroic mirror before the camera installed on the end bench.
After solving this issue and installing the AOM we will start to look for the frequency shift needed to have both the IR and the green resonant on the fundamental mode.

Then we powered up the AOM, there was the other orders appearing, the one close to the 0 order is the 1st order which we need. We put a PD at the 1st order, connected it to the oscilloscope and started to turn the AOM to have maximum power in the 1st order. After finding the best position of the AOM, we tried to change the RF power we were sending to increase the power more. But when we tried to do this, we found out the power fluctuation is about 40-50% percent, it was much more than the tolerance.
We put PD both in the infrared path and the green path to check the fluctuation of the green is coming from the infrared or not. Then we found out since the PD aperture is quite small, so it will increase the degree of uncertainty we saw. So we changed the PD to the powermeter which has larger aperture. This time the infrared is quite stable and the green is fluctuating but much less than before. We guessed that the fluctuation of the green is coming from the changing of the SHG cavity alignment when people moving around or touching the bench. We are going the install the MZ and stabilize the green.
Then other thing we checked is that the SHG loop gain we are using now is 50. But with this gain,if we changed the power sent to the PZT, the cavity cannot pull the error signal back to zero when the cavity is locked. If we increase the gain, until it reaches 10000, there will be no oscillation appear, the cavity will be locked more stable, but one problem is that it will be quite easy also lock on the other mode. We will try to find a the optimal gain considering also the filter cavity.

- In order to increase the pump power safely, we cover the laser path with aluminum black walls and black paper.
- We also cover the probe part in order to stop air flow from hepa filters which is a possible noise source. Kuroki-san helped me on this tasks.
- I put a small translation stage below the half ball in the imaging unit in order to adjust better the alignment and I tried many times the alignment until I found a good signal for the surface reference sample, comparable with the one of last year. Also the bulk reference sample gives almost the same signal as last year.
- The parts for the 1310nm probe laser were delivered, I glued the golden prism mirror on a half inch post, I'm waiting for the glue to cure under the neon lamps



The PDH filter cavity signal has been calibrated injecting a line at 28 kHz (above the ugf ~ 10 kHz of the loop) on the “ramp” input of the electronic servo. The ramp input is summed to the PZT correction signal.
The amplitude of the 28 kHz line in Hz is obtained using the formula:
S_Hz = V_RMS (V) * sqrt(2)*100*2e6 Hz/V = 1.25e-6*sqrt(2)*100*2e6 = 353 Hz
Where V_RMS is the line amplitude measured by the Agilent spectrum analyser. The factor sqrt(2) is obtained to pass from the V_RMS to the line amplitude (the factor has been also experimentally verified looking the same line with the spectrum analyser and the oscilloscope).
The factor 100 is the reduction of the PZT_moni output . 2e6 Hz/V is the calibration of the PZT after the SHG.
Measuring the line at 28 kHz in the error signal and compensating for the cavity frequency pole is it possible to find the calibration factor K in V/Hz. The formula used is :
S_V = K(V/Hz)* S_Hz /sqrt(1+ (f/f_0)^2)
where f_0 = 1.5 kHz and S_V = sqrt(2)*38.9e-3 V
--> K = 2.9e-3 V/Hz
which seems to be in agreement with the calibration obtained looking the PDH signal when the cavity is freely swinging. In that case we see a peak-to-peak of the PDH of ~ 4 V for 1.5 kHz of the cavity line which correspond to a K = 2.7e-3 V/Hz. Note that when the cavity is freely swining we have also rining effects which can perturb this measurement.
We have also checked that reducing the frequency of the line sent to the PZT (with the same amplitude) to 14 kHz, the amplitude of the line of the error signal is multiplied by 2, as expected given the cavity pole. A more quantitative analysis (fully taking in account the effect of the loop) is necessary to check the position of the cavity pole.
Another test was to increase the amplitude of the line by a factor 10, thus having a 29 kHz line with amplitude of 3 kHz (two times the cavity width of the cavity). The cavity stays locked and the calibration factor measured is the same with the one measured with the line with an amplitude of 300 Hz. Increasing further the amplitude of the 28 kHz line to ~ 7 kHz (4 times the cavity linewidth) makes the lock more fragile, and sometimes the cavity unlocks. Moreover, an oscillation with a frequency of ~ 1 Hz appears in the error signal (but it is not accompanied with a similar oscillation in the transmitted power).